OpenStudy (anonymous):

Find the measure of angle A if a = 34, b = 28 and c = 15

OpenStudy (anonymous):

I have this but I dont know what should I do next. 34^2= 28^2+15^2 -2(28*15) COS A 6.84 = Cos A

OpenStudy (anonymous):

cosine A = (-a² + b² + c²)/ABC A = arccos((-a² + b² + c²)/abc) = arccos(-34² + 28² + 15²)/14280 = arccos(-147/14280) = 90.59° :)

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

your very welcome ! :)